OpenStudy (anonymous):

3(6z-1)-3(z+7)=13(z+1)

OpenStudy (precal):

3(6z-1)-3(z+7)=13(z+1) 18z-3-3z-21=13z+13 15z-24=13z+13 2z=37 z=37/2